OpenStudy (anonymous):

In the proportion 3:7 = 6:14, name the means and the extremes. Means are 3 and 7 extremes are 6 and 14 Means are 6 and 14 extremes are 3 and 7 Means are 7 and 6 extremes are 3 and 14 Means are 3 and 14 extremes are 7 and 6

OpenStudy (anonymous):

means are 7 and 6, extremes are 3 and 14

OpenStudy (anonymous):

which of the following represents a true proportion. 2/3= 4/6, 7:2= 8:5, 6:5=2:10, 1/6= 6/1.?

OpenStudy (anonymous):

and how are you doing this?

OpenStudy (anonymous):

2/3= 4/6

OpenStudy (anonymous):

product of means = product of extremes

OpenStudy (anonymous):

which is false for all the other proportions